OpenStudy (zaxoanl):

A 0.18-mole sample of nitrogen gas occupies a volume of 2.95L . What is the volume of 0.37mol of nitrogen gas under the same conditions?

OpenStudy (zaxoanl):

i got 1.435 why is it wrong please help .18*2.95 / .37

OpenStudy (abhisar):

Hello @zaxoanl ! \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (abhisar):

V\(\propto\)n So, \(\sf\large \frac{V_1}{V_2}=\frac{n_1}{n_2}\\~\\~\\-->\large \frac{2.95}{V}=\frac{0.18}{0.37}\)

OpenStudy (abhisar):

Now find the value of V

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.